Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name Jixie
Log entry time 06:02:35 on March10,2012
Entry number 365317
keyword=Wrong momentum control for both arm from Swing shift
During optics data taking I found serious problem in the optics data. Then I looking into the issue and found that Q1 is at a wrong current. After looked deep into it I found that the m0omentum setting is WRONG!!! It should be 'g2p|gEp_1' but it was 'no_septum'!!!
I check the Q1 archive and found that Q1 was wrong since 16:00! Therefore all data we took before 6:00 AM today is not useful at all.
I highly recommend to remove 'no_septum' from the contron such that this mistake will not happen again.
FIGURE 1
FIGURE 2